# Conformalized Credal Regions for Classification with Ambiguous Ground Truth

### v1

@@ 4. Conformal Plausibility Regions as Credal Regions
The plausibility region $c(X_{n+1})$ derived in (1) is convex and closed. [^6] The topology in which $c(X_{n+1})$ is closed is specified in the proof of the statement.
@@
We first show that $c(X_{n+1})$ is convex. Pick $\lambda^{1},\lambda^{2}\in c(X_{n+1})$, $\lambda^{1}\neq\lambda^{2}$, and $\beta\in[0,1]$. Then,
$e(X_{n+1},\beta\lambda^{1}+(1-\beta)\lambda^{2})$ $=\sum_{k=1}^{K}(\beta\lambda^{1}_{k}+(1-\beta)\lambda^{2}_{k})E(X_{n+1},k)$
$=\beta\sum_{k=1}^{K}\lambda^{1}_{k}E(X_{n+1},k)+(1-\beta)\sum_{k=1}^{K}\lambda^{2}_{k}E(X_{n+1},k)$
@@
Hence, $\beta\lambda^{1}+(1-\beta)\lambda^{2}\in c(X_{n+1})$, which proves convexity.
Let us then turn our attention to sequential closure. We begin by showing that $e(X_{n+1},\cdot)$ is a continuous function. Pick $(\lambda^{m})\subseteq c(X_{n+1})$ such that $\lambda^{m}\rightarrow\lambda^{\star}$. That is, for all $\gamma>0$, there exists $M\in\mathbb{N}$ such that $|\lambda^{\star}_{k}-\lambda^{m}_{k}|<\gamma$, for all $m\geq M$ and all $k\in\{1,\ldots,K\}$. Now, to show continuity, we prove that, for all $\epsilon>0$, there exists $\delta_{\epsilon}\coloneqq\epsilon/\sum_{k=1}^{K}E(X_{n+1},k)>0$ such that $|\lambda^{\star}_{k}-\lambda^{m}_{k}|<\delta_{\epsilon}$ implies that $|e(X_{n+1},\lambda^{\star})-e(X_{n+1},\lambda^{m})|<\epsilon$. Indeed,
$\left|e(X_{n+1},\lambda^{\star})-e(X_{n+1},\lambda^{m})\right|$ $=\left|\sum_{k=1}^{K}(\lambda^{\star}_{k}-\lambda^{m}_{k})E(X_{n+1},k)\right|$
$\leq\sum_{k=1}^{K}\left|\lambda^{\star}_{k}-\lambda^{m}_{k}\right|E(X_{n+1},k)$
@@
This proves that $e(X_{n+1},\cdot)$ is continuous. Then, after noting that $e(X_{n+1},\lambda)\geq\tau$, for all $\lambda\in c(X_{n+1})$ – here the weak inequality plays a crucial role in showing sequential closure – we can conclude that $c(X_{n+1})$ is indeed sequentially closed. ∎
Figure 1. Suppose we are in a $3$-class classification setting, so $\mathcal{Y}=\{1,2,3\}$. Then, any probability measure $P$ on $\mathcal{Y}$ can be seen as a probability vector. For example, suppose $P(\{1\})=0.6$, $P(\{2\})=0.3$, and $P(\{3\})=0.1$. We have that $P\equiv(0.6,0.3,0.1)^{\top}$. Since its elements are positive and sum up to $1$, probability vector $P$ belongs to the unit simplex $\Delta^{2}$ in $\mathbb{R}^{3}$, the purple (2D) triangle in the figure. Then, plausibility region $c(X_{n+1})$ is a closed and convex body in $\Delta^{2}$, such as the depicted orange (2D) heptagon (we depict it with a solid black border to highlight the fact that it is indeed closed).
As we can see, then, every $\lambda\in c(X_{n+1})$ uniquely identifies a Categorical distribution $\textup{Cat}(\lambda)$ parameterized by $\lambda$ itself. A visual representation for $c(X_{n+1})$ is given in Figure 1. A consequence of Proposition 4 is that for all $k\in\mathcal{Y}$, we can find $\underline{\lambda}_{k},\overline{\lambda}_{k}\in[0,1]$, $\underline{\lambda}_{k}\leq\overline{\lambda}_{k}$, such that $\lambda_{k}\in[\underline{\lambda}_{k},\overline{\lambda}_{k}]$, for all $\lambda\in c(X_{n+1})$. [^7] As a consequence, a greedy algorithm to approximate the values of $\underline{\lambda}_{k},\overline{\lambda}_{k}$, for all $k\in\mathcal{Y}$, is easy to design, e.g. based on random samples from the uniform $\mathrm{Unif}(c(X_{n+1}))$ on the plausibility region $c(X_{n+1})$. The following result relates $c(X_{n+1})$ to a credal region, that is, a closed and convex family of probabilities [34]. It is an immediate consequence of Proposition 4 and of the definition of $c(X_{n+1})$ in (1).
@@
The plausibility region $c(X_{n+1})$ derived in (1) is equivalent to the credal region $\mathcal{P}\coloneqq\{\textup{Cat}(\lambda):\lambda\in c(X_{n+1})\}$.
@@
Let $\alpha$ be the same significance level selected in (1). Then,
@@
where $\mathbb{P}$ depends on the statistical model relating expert opinions to plausibilities as outlined in Section 2 and [53, Equation (6)].
@@
Immediate from Corollary 4.1 and [53, Equation (13)]. ∎
4.1. Significance for IPML
@@
In this section, we argue that Proposition 5 is extremely important for the field of Imprecise Probabilistic Machine Learning (IPML) [6, 30, 56]. We will further show (in the next section) that making this connection to imprecise probability allows to derive more efficient predictive sets from these credal regions, compared to $\Psi(c(X_{n+1}))$ in (2).
@@
In IPML, scholars tend to adopt either of the following two approaches. One approach is the robust statistician one, also called Huberian [28]. This is more frequentist in nature, hence the true data generating process – $\textup{Cat}(\lambda_{n+1})$ in our case – is a well-defined concept. Those who take this approach proceed in two different ways: either they assume that $\textup{Cat}(\lambda_{n+1})\in\mathcal{P}$ (in which case a result like Proposition 5 is very interesting, as it allows to forego such requirement), or they verify that a calibration à la Proposition 5 holds for the methodology they propose [2, 26, 42, 37, 16]. In particular, in the case of transductive conformal prediction, [39] shows that the upper envelope $\overline{\Pi}$ of the credal region $\mathcal{M}(\overline{\Pi})$ induced by conformal transducer $\pi$ as we discussed in Section 1, is the minimal outer consonant approximation of the true data generating process. This means that $\overline{\Pi}$ is the narrowest upper bound for the true distribution, that also satisfies the consonance assumption. We can then conclude – although it is not explicitly shown in their work – that credal region $\mathcal{M}(\overline{\Pi})$ contains the true data generating process. This is important in Cella and Martin’s framework because it is linked to the concept of type-2 validity: since the upper probability $\overline{\Pi}$ provides control on erroneous predictions uniformly [15, Definition 3], so does the true probability if the credal set is well-calibrated. Of course, we cannot directly use their result in our context because we do not have access to crisp labels. In a sense, then, in Proposition 5 we prove (a probabilistic version of) this property for the ambiguous ground truth case in classification problems.
@@
Another approach is the deFinettian or Walleyan one [20, 21, 54], in which it is posited that “probability does not exist”, and what we call probability measure is just a way of capturing the subjective assessments of the likelihood of events to obtain, according to the agent. In light of this, the concept of “true data generating process” is absent, and so it does not make sense to look for probabilistic guarantees à la Proposition 5.
4.2. Relation to [14, 15]
@@
Given calibration data with crisp labels in $\mathcal{Y}$ as outlined in the beginning of Section 2, [14, 15] consider the conformal transducer $\pi:\mathcal{Y}\rightarrow[0,1]$, which assigns a $p$-value to each possible label in $\mathcal{Y}$ [14, Definition 2]. It uniquely identifies a credal region, assuming consonance. Here, $\pi$ is typically interpreted as telling us how “in line” the pair $(x_{n+1},\tilde{y})$ is with the previously observed data $(\mathbf{x^{n}},\mathbf{y^{n}})$. A value closer to $1$ means that seeing $(X_{n+1},Y_{n+1})=(x_{n+1},\tilde{y})$ would align with the preceding observations, and vice versa for values closer to 00. Once $\pi$ is obtained, consonance posits that there exists at least one label $\tilde{y}$ such that $\pi(\tilde{y})=1$. Loosely, this means that the label $\tilde{y}$ that makes pair $(x_{n+1},\tilde{y})$ “the most conformal to” the observations $(\mathbf{x^{n}},\mathbf{y^{n}})$, is required to have the highest possible value $1$. This can be obtained artificially by setting $\pi$ for $\tilde{y}\in\text{argmax}_{y}\pi(y)$ to $1$ [14, Section 7]. Then, an upper probability (that is, the upper envelope of a credal region) $\overline{\Pi}$ is derived for $\pi$ by letting $\overline{\Pi}(A)=\sup_{y\in A}\pi(y)$, for all subsets $A$ of the label space $\mathcal{Y}$. In turn, a credal region is defined as $\mathcal{M}(\overline{\Pi})=\{P:P(A)\leq\overline{\Pi}(A)\}$, for all $A\subseteq\mathcal{Y}$. That is, $\mathcal{M}(\overline{\Pi})$ contains all the probabilities on $\mathcal{Y}$ that are set-wise dominated by $\overline{\Pi}$ where $P(A)\equiv P(Y_{n+1}\in A)$.
@@
In comparison to [14, 15], we avoid the consonance assumption. In [14, 15], this assumptions is primarily required to bridge the so-called possibilistic approach to imprecise probabilities [3, Chapter 4] with conformal prediction. Moreover, we do not explicitly construct the conformal transducer, also due to the fact that we work with probability vectors instead of “precise” (that is, one-hot) labels. In this sense, our work subsumes and extends [14, 15]. Furthermore, our credal region $\mathcal{P}$ establishes a coverage guarantee on the true data generating process being in $\mathcal{P}$, and it also enjoys (a version of) type-2 validity, a notion important for uncertainty quantification, introduced for the first time in [15, Definition 2]. Let us be more formal about it.
@@
The credal region $\mathcal{P}$ in Corollary 4.1 is $[\delta/(1-\alpha)]$-type-2 valid, where $\alpha$ is the quantity chosen in (1). That is,
$\mathbb{P}[\overline{P}(A)\leq\delta\text{, }Y_{n+1}\in A]\leq\frac{\delta}{1-\alpha},$
@@
Pick any $\delta\in[0,1]$, any $n\in\mathbb{N}$, and any $A\subseteq\mathcal{Y}$. Notice that $\mathbb{P}[\overline{P}(A)\leq\delta\text{, }Y_{n+1}\in A]\leq\mathbb{P}[Y_{n+1}\in A]$. If $\textup{Cat}(\lambda_{n+1})\in\mathcal{P}$, which happens with $\mathbb{P}$-probability $(1-\alpha)$ as shown in Proposition 5, then $\mathbb{P}[Y_{n+1}\in A]\leq\overline{P}(A)$, which in turn is smaller than $\delta$ by choice. Hence, the claim follows by dividing $\delta$ by $(1-\alpha)$. ∎
In words, this means that the (true/aggregated) probability of a set $A$ having small upper probability, and yet containing the true output $Y_{n+1}$ for a given new input $X_{n+1}$, is controllably small, and it depends on the parameters $\alpha$ and $\delta$ chosen by the user.
5. Improving over Conformal Prediction Sets
In this section, we show how prediction set $\Psi(c(X_{n+1}))$ can be improved by using the IHDS of the lower probability $\underline{P}$ of credal region $\mathcal{P}$. In addition, we show how $\mathcal{P}$ can be used to quantify and disentangle aleatoric and epistemic uncertainties. Specifically, let us denote by $\underline{\lambda}$ the vector whose $k$-th entry is $\underline{\lambda}_{k}=\underline{P}(\{k\})$, and by $\overline{\lambda}$ the vector whose $k$-th entry is $\overline{\lambda}_{k}=\overline{P}(\{k\})$. Then, we can rewrite (3) as

### v2

@@ 3 Conformal Plausibility Regions as Credal Regions
The plausibility region $c(X_{n+1})$ derived in equation 1 is convex and closed. [^6] The topology in which $c(X_{n+1})$ is closed is specified in the proof of the statement.
@@
We first show that $c(X_{n+1})$ is convex. Pick $\lambda^{(1)},\lambda^{(2)}\in c(X_{n+1})$, $\lambda^{(1)}\neq\lambda^{(2)}$, and $\beta\in[0,1]$. Then,
$e(X_{n+1},\beta\lambda^{(1)}+(1-\beta)\lambda^{(2)})$ $=\sum_{k=1}^{K}(\beta\lambda^{(1)}_{k}+(1-\beta)\lambda^{(2)}_{k})E(X_{n+1},k)$
$=\beta\sum_{k=1}^{K}\lambda^{(1)}_{k}E(X_{n+1},k)+(1-\beta)\sum_{k=1}^{K}\lambda^{(2)}_{k}E(X_{n+1},k)$
@@
Hence, $\beta\lambda^{(1)}+(1-\beta)\lambda^{(2)}\in c(X_{n+1})$, which proves convexity.
Let us then turn our attention to closure. We begin by showing that $e(X_{n+1},\cdot)$ is a continuous function. Pick a converging sequence $(\lambda^{(m)})\subseteq c(X_{n+1})$. This means that there is a probability vector $\lambda^{\star}$ such that $\lambda^{(m)}\rightarrow\lambda^{\star}$. That is, for all $\gamma>0$, there exists $M\in\mathbb{N}$ such that $|\lambda^{\star}_{k}-\lambda^{(m)}_{k}|<\gamma$, for all $m\geq M$ and all $k\in\{1,\ldots,K\}$. Now, to show continuity, we prove that, for all $\epsilon>0$, there exists $\delta_{\epsilon}\coloneqq\epsilon/\sum_{k=1}^{K}E(X_{n+1},k)>0$ such that $|\lambda^{\star}_{k}-\lambda^{(m)}_{k}|<\delta_{\epsilon}$ implies that $|e(X_{n+1},\lambda^{\star})-e(X_{n+1},\lambda^{(m)})|<\epsilon$. Indeed,
$\left|e(X_{n+1},\lambda^{\star})-e(X_{n+1},\lambda^{(m)})\right|$ $=\left|\sum_{k=1}^{K}(\lambda^{\star}_{k}-\lambda^{(m)}_{k})E(X_{n+1},k)\right|$
$\leq\sum_{k=1}^{K}\left|\lambda^{\star}_{k}-\lambda^{(m)}_{k}\right|E(X_{n+1},k)$
@@
This proves that $e(X_{n+1},\cdot)$ is continuous. Then, after noting that $e(X_{n+1},\lambda)\geq\tau$, for all $\lambda\in c(X_{n+1})$ – here the weak inequality plays a crucial role in showing closure – we can conclude that $e(X_{n+1},\lambda^{\star})\geq\tau$, and so $\lambda^{\star}\in c(X_{n+1})$, proving that $c(X_{n+1})$ is indeed closed. ∎
Figure 1: Suppose we are in a $3$-class classification setting, so $\mathcal{Y}=\{1,2,3\}$. Then, any probability measure $P$ on $\mathcal{Y}$ can be seen as a probability vector. For example, suppose $P(\{1\})=0.6$, $P(\{2\})=0.3$, and $P(\{3\})=0.1$. We have that $P\equiv(0.6,0.3,0.1)^{\top}$. Since its elements are positive and sum up to $1$, probability vector $P$ belongs to the unit simplex $\Delta^{2}$ in $\mathbb{R}^{3}$, the purple (2D) triangle in the figure. Then, plausibility region $c(X_{n+1})$ is a closed and convex body in $\Delta^{2}$, such as the depicted orange (2D) heptagon (we depict it with a solid black border to highlight the fact that it is indeed closed).
As we can see, then, every $\lambda\in c(X_{n+1})$ uniquely identifies a Categorical distribution $\textup{Cat}(\lambda)$ parameterized by $\lambda$ itself. A visual representation for $c(X_{n+1})$ is given in Figure 1. A consequence of Proposition 4 is that for all $k\in\mathcal{Y}$, we can find $\underline{\lambda}_{k},\overline{\lambda}_{k}\in[0,1]$, $\underline{\lambda}_{k}\leq\overline{\lambda}_{k}$, such that $\lambda_{k}\in[\underline{\lambda}_{k},\overline{\lambda}_{k}]$, for all $\lambda\in c(X_{n+1})$. [^7] As a consequence, a greedy algorithm to approximate the values of $\underline{\lambda}_{k},\overline{\lambda}_{k}$, for all $k\in\mathcal{Y}$, is easy to design, e.g. based on random samples from the uniform $\mathrm{Unif}(c(X_{n+1}))$ on the plausibility region $c(X_{n+1})$. The following result relates $c(X_{n+1})$ to a credal region, that is, a closed and convex family of probabilities (Levi, 1980). It is an immediate consequence of Proposition 4 and of the definition of $c(X_{n+1})$ in equation 1.
@@
The plausibility region $c(X_{n+1})$ derived in equation 1 is equivalent to the credal region $\mathcal{P}\coloneqq\{\textup{Cat}(\lambda):\lambda\in c(X_{n+1})\}$.
@@
Let $\alpha$ be the same significance level selected in equation 1. Then,
@@
where $\mathbb{P}$ depends on the statistical model relating expert opinions to plausibilities as outlined in Section 2.1 and (Stutz et al., 2023c, Equation (6)).
@@
Immediate from Corollary 4.1 and (Stutz et al., 2023c, Equation (13)). ∎
3.1 Significance for IPML
@@
In this section, we argue that Proposition 5 is extremely important for the field of Imprecise Probabilistic Machine Learning (IPML) (Caprio et al., 2024a; Hüllermeier & Waegeman, 2021; Zaffalon, 2002). We will further show (in the next section) that making this connection to imprecise probability allows us to derive more efficient predictive sets from these credal regions, compared to $\Psi(c(X_{n+1}))$ in equation 2.
@@
In IPML, scholars tend to adopt either of the following two approaches. One approach is the robust statistician one, also called Huberian (Huber & Ronchetti, 2009). This is more frequentist in nature, hence the true data generating process – $\textup{Cat}(\lambda_{n+1})$ in our case – is a well-defined concept. Those who take this approach proceed in two different ways: either they assume that $\textup{Cat}(\lambda_{n+1})\in\mathcal{P}$ (in which case a result like Proposition 5 is very interesting, as it allows to forego such requirement), or they verify that a calibration à la Proposition 5 holds for the methodology they propose (Acharya et al., 2015; Gao et al., 2018; Mortier et al., 2023; Liu & Briol, 2024; Chau et al., 2024). In particular, in the case of transductive conformal prediction, Martin (2023) shows that the upper envelope $\overline{\Pi}$ of the credal region $\mathcal{M}(\overline{\Pi})$ induced by conformal transducer $\pi$ as we discussed in Section 1, is the minimal outer consonant approximation of the true data generating process. This means that $\overline{\Pi}$ is the narrowest upper bound for the true distribution, that also satisfies the consonance assumption. We can then conclude – although it is not explicitly shown in their work – that credal region $\mathcal{M}(\overline{\Pi})$ contains the true data generating process. This is important in Cella and Martin’s framework because it is linked to the concept of type-2 validity: since the upper probability $\overline{\Pi}$ provides control on erroneous predictions uniformly (Cella & Martin, 2022b, Definition 3), so does the true probability if the credal region is well-calibrated. Of course, we cannot directly use their result in our context because we do not have access to crisp labels. In a sense, then, in Proposition 5 we prove (a probabilistic version of) this property for the ambiguous ground truth case in classification problems.
@@
Another approach is the deFinettian or Walleyan one (de Finetti, 1974; 1975; Walley, 1991), in which it is posited that “probability does not exist”, and what we call probability measure is just a way of capturing the subjective assessments of the likelihood of events to obtain, according to the agent. In light of this, the concept of “true data generating process” is absent, and so it does not make sense to look for probabilistic guarantees à la Proposition 5.
3.2 Relation to Cella & Martin (2022a; b)
@@
Given calibration data with crisp labels in $\mathcal{Y}$ as outlined in the beginning of Section 2.1, Cella & Martin (2022a; b) consider the conformal transducer $\pi:\mathcal{Y}\rightarrow[0,1]$, which assigns a $p$-value to each possible label in $\mathcal{Y}$ (Cella & Martin, 2022a, Definition 2). It uniquely identifies a credal region, assuming consonance. Here, $\pi$ is typically interpreted as telling us how “in line” the pair $(x_{n+1},\tilde{y})$ is with the previously observed data $(\mathbf{x^{n}},\mathbf{y^{n}})$. A value closer to $1$ means that seeing $(X_{n+1},Y_{n+1})=(x_{n+1},\tilde{y})$ would align with the preceding observations, and vice versa for values closer to 00. Once $\pi$ is obtained, consonance posits that there exists at least one label $\tilde{y}$ such that $\pi(\tilde{y})=1$. Loosely, this means that the label $\tilde{y}$ that makes pair $(x_{n+1},\tilde{y})$ “the most conformal to” the observations $(\mathbf{x^{n}},\mathbf{y^{n}})$, is required to have the highest possible value $1$. This can be obtained artificially by setting $\pi$ for $\tilde{y}\in\operatorname*{arg\,max}_{y}\pi(y)$ to $1$ (Cella & Martin, 2022a, Section 7). Then, an upper probability (that is, the upper envelope of a credal region) $\overline{\Pi}$ is derived for $\pi$ by letting $\overline{\Pi}(A)=\sup_{y\in A}\pi(y)$, for all subsets $A$ of the label space $\mathcal{Y}$. In turn, a credal region is defined as $\mathcal{M}(\overline{\Pi})=\{P:P(A)\leq\overline{\Pi}(A)\}$, for all $A\subseteq\mathcal{Y}$. That is, $\mathcal{M}(\overline{\Pi})$ contains all the probabilities on $\mathcal{Y}$ that are set-wise dominated by $\overline{\Pi}$ where $P(A)\equiv P(Y_{n+1}\in A)$.
@@
In comparison to Cella & Martin (2022a; b), we avoid the consonance assumption. In Cella & Martin (2022a; b), this assumptions is primarily required to bridge the so-called possibilistic approach to imprecise probabilities (Augustin et al., 2014, Chapter 4) with conformal prediction. Moreover, we do not explicitly construct the conformal transducer, also due to the fact that we work with probability vectors instead of “precise” (that is, one-hot) labels. In this sense, our work subsumes and extends Cella & Martin (2022a; b). Furthermore, our credal region $\mathcal{P}$ establishes a coverage guarantee on the true data generating process being in $\mathcal{P}$, and it also enjoys (a version of) type-2 validity, a notion important for uncertainty quantification, introduced for the first time in (Cella & Martin, 2022b, Definition 2). Let us be more formal about it.
@@
The credal region $\mathcal{P}$ in Corollary 4.1 is $[\delta/(1-\alpha)]$-type-2 valid, where $\alpha$ is the quantity chosen in equation 1. That is,
$\mathbb{P}[\overline{P}(A)\leq\delta\text{, }Y_{n+1}\in A]\leq\frac{\delta}{1-\alpha},$ (4)
@@
Pick any $\delta\in[0,1]$, any $n\in\mathbb{N}$, and any $A\subseteq\mathcal{Y}$. Notice that $\mathbb{P}[\overline{P}(A)\leq\delta\text{, }Y_{n+1}\in A]\leq\mathbb{P}[Y_{n+1}\in A]$. Now, if $\textup{Cat}(\lambda_{n+1})\in\mathcal{P}$, then $\mathbb{P}[Y_{n+1}\in A]\leq\overline{P}(A)$. In addition, given our assumption on $\delta$, we know that $\overline{P}(A)\leq\delta$. Hence, we can conclude that if $\textup{Cat}(\lambda_{n+1})\in\mathcal{P}$, then $\mathbb{P}[\overline{P}(A)\leq\delta\text{, }Y_{n+1}\in A]\leq\delta$. But $\textup{Cat}(\lambda_{n+1})\in\mathcal{P}$ happens with $\mathbb{P}$-probability $(1-\alpha)$ as shown in Proposition 5, and so we have that $\mathbb{P}[\overline{P}(A)\leq\delta\text{, }Y_{n+1}\in A]\leq\delta/[1-\alpha]$. ∎
In words, this means that the (true/aggregated) probability of a set $A$ having small upper probability, and yet containing the true output $Y_{n+1}$ for a given new input $X_{n+1}$, is controllably small, and it depends on the parameters $\alpha$ and $\delta$ chosen by the user. When giving the bound, we need to divide by $1-\alpha$ because we need to take into account what is the $\mathbb{P}$-probability that the true distribution $\textup{Cat}(\lambda_{n+1})$ belongs to the credal region $\mathcal{P}$. Notice that in equation 4 we find a slightly looser bound than the one we would derive if we were not dealing with ambiguous ground truth, i.e. if we knew that $\mathbb{P}[\textup{Cat}(\lambda_{n+1})\in\mathcal{P}]=1$. To see that this is indeed the case, notice that if we pick $\delta=\alpha=0.05$, then the bound in equation 4 is $\approx 0.526$, slightly larger than the one we would have ($0.5$) if we were not facing ambiguity.
4 Improving over Conformal Prediction Sets
In this section, we show how prediction set $\Psi(c(X_{n+1}))$ can be improved by using the IHDS of the lower probability $\underline{P}$ of credal region $\mathcal{P}$. In addition, we show how $\mathcal{P}$ can be used to quantify and disentangle aleatoric and epistemic uncertainties. Specifically, let us denote by $\underline{\lambda}$ the vector whose $k$-th entry is $\underline{\lambda}_{k}=\underline{P}(\{k\})$, and by $\overline{\lambda}$ the vector whose $k$-th entry is $\overline{\lambda}_{k}=\overline{P}(\{k\})$. Then, we can rewrite equation 3 as